Phillies' second baseman, Chase Utley, has a batting average of .289. Assume that's the probability of him getting a hit in any
single at bat. If he has three at bats, find the probability that :
a. He gets three hits.


b. He gets no hits.


c. He gets at least one hit.

Answer by ewatrrr(24785)

p(hit) = .289
a) P(x = 3) = (.289)^3
b) P(x = 0) = (.711)^3
c. P(at least one) = 1 - P(none) = 1 - (.771)^3